# Coexistence of Chiral Majorana Edge States and Bogoliubov Fermi Surfaces in Two-Dimensional Nonsymmorphic Dirac Semimetal/Superconductor Heterostructures

### v1

@@ Abstract
where $\epsilon_{\pm}(\boldsymbol{k})=\epsilon_{0}(\boldsymbol{k})-\mu\pm\sqrt{\epsilon_{1}(\boldsymbol{k})^{2}+\Lambda^{2}(\boldsymbol{k})}$. It is evident that $\eta\neq 0$ (so $\epsilon_{1}(\boldsymbol{k})$ does not vanish identically) and $|M_{z}|>|\Delta_{s}|$ are two necessary conditions for $P(\boldsymbol{k})$ to have zeros in this system. In Fig.4, we illustrate two representative configurations of BFSs by plotting the contours where $P(\boldsymbol{k})$ equals zero. We have verified that the sign of $P(\boldsymbol{k})$ indeed changes across these BFSs, confirming their topological protection.
Figure 4: Two representative configurations of BFSs (red lines) in the Brillouin zone. Panel (a) corresponds to parameters $(\mu,\eta)=(4,1)$, while panel (b) corresponds to $(\mu,\eta)=(0,0.3)$. The common parameters for both figures are $t=1$, $\lambda=0.5$, $M_{z}=0.5$ and $\Delta_{s}=0.2$.
The impact of BFSs on thermal Hall effects.—A defining characteristic of CTSCs is the quantized thermal Hall effect in the low-temperature limit. The quantized coefficient in this effect is directly proportional to the Chern number, expressed as $\kappa_{xy}=C\kappa_{0}$, where $\kappa_{0}=\pi k_{B}^{2}T/12\hbar$[72, 73]. When a bulk energy gap exists, low-energy excitations are confined to the boundary, and the quantized effect can be derived using conformal field theory. This theory elucidates that each branch of the chiral Majorana edge state contributes equally to the effect[9]. However, in the presence of BFSs, low-energy excitations also occur within the bulk. These low-energy Bogoliubov quasiparticles near the BFSs experience a nonzero Berry curvature, resulting in a nonzero contribution to the thermal Hall effect that is generally non-quantized[74]. To investigate the influence of BFSs on the thermal Hall effect, we utilize the formula for the thermal Hall coefficient developed in Refs.[75, 76], which is given by
[…]
where $f(E)=1/[\exp(E/k_{B}T)+1]$ is the Fermi-Dirac distribution function, $k_{B}$ is the Boltzmann constant, $T$ is the temperature, and
$\sigma(E)=\sum_{n}\int_{E_{k}<E}\frac{d^{2}k}{(2\pi)^{2}}\Omega^{(n)}(\boldsymbol{k}).$ (13)
@@
Figure 5: Evolution of the thermal Hall coefficient with respect to $\eta$. The common parameters are $t=1$, $\lambda=0.5$, $M_{z}=0.5$ and $\Delta_{s}=0.2$. (a) $\mu=-4$, (b) $\mu=0$, (c) $\mu=1$, (d) $\mu=4$. The critical value of $\eta$ for the onset of BFSs is approximately equal to 0.11 in (a), 0.19 in (b), 0.27 in (c) and 0.72 in (d). The Chern number before the onset of BFS is equal to $2$ in (a), $-4$ in (b), 00 in (c) and $2$ in (d).
@@
In Fig.5, we present the evolution of the $\kappa_{xy}$ with respect to $\eta$. In Fig.5(a), 5(b), 5(c) and 5(d), the Chern number is respectively $2$, $-4$, 00, and $2$ before $\eta$ surpasses the critical value, denoted as $\eta_{c}$, where BFSs emerge. Notably, $\eta_{c}$ is dependent on other parameters, as can be inferred from Eq.(11). The results depicted in Fig.5 demonstrate quantized behavior in the low-temperature limit, provided that the bulk energy gap remains open. As anticipated, once $\eta$ exceeds $\eta_{c}$, the quantized behavior diminishes due to the presence of BFSs. Intriguingly, we observe that the change in $\kappa_{xy}$ due to BFSs is most pronounced in Fig.5(a). For the largest value of $\eta$ considered, $\kappa_{xy}$ in Fig.5(a) deviates significantly from its original quantized value, while the results in Fig.5(b) and Fig.5(d) show a much smaller derivation from their original quantized values. The remarkable change observed in Fig.5(a) can be attributed to the fact that the spectrum crossing of chiral Majorana edge states occurs at $k_{x}=0$ or $k_{y}=0$ for this particular case,. As previously discussed, the presence of BFSs will cause these chiral Majorana edge states to disappear.
@@
Discussions and conclusions.—We demonstrated that applying a Zeeman field perpendicular to the spin plane of the inversion-symmetric SOC can transform the nonsymmorphic DSM into a Chern insulator. Furthermore, we showed that incorporating $s$-wave superconductivity into this system can lead to the realization of fully-gapped CTSCs with large Chern numbers, as well as gapless superconducting phases where BFSs coexist with chiral Majorana edge states. Additionally, we found that the presence of BFSs can have a significant impact on the thermal Hall effect. Our study underscores the rich interplay between Zeeman fields, superconductivity, and nonsymmorphic Dirac band structures, which can give rise to a diverse array of phases exhibiting intriguing properties.
@@
Acknowledgements.— This work is supported by the National Natural Science Foundation of China (Grant No.12174455) and the Natural Science Foundation of Guangdong Province (Grant No. 2021B1515020026).

### v2

@@ Abstract
where $\epsilon_{\pm}(\boldsymbol{k})=\epsilon_{0}(\boldsymbol{k})-\mu\pm\sqrt{\epsilon_{1}^{2}(\boldsymbol{k})+\Lambda^{2}(\boldsymbol{k})}$. It is evident that $\eta\neq 0$ (ensuring $\epsilon_{1}(\boldsymbol{k})$ does not vanish identically) and $|M_{z}|>|\Delta_{s}|$ are two necessary conditions for $P(\boldsymbol{k})$ to have zeros in this system. In Fig.4, we illustrate two representative configurations of BFSs by plotting the contours where $P(\boldsymbol{k})$ equals zero. We have verified that the sign of $P(\boldsymbol{k})$ indeed changes across these BFSs, confirming their topological protection.
Figure 4: Two representative configurations of BFSs (red lines) in the Brillouin zone. Panel (a) corresponds to parameters $(\mu,\eta)=(4,1)$, while panel (b) corresponds to $(\mu,\eta)=(0,0.3)$. The values of the shared parameters are: $t=1$, $\lambda_{\rm so}=0.5$, $M_{z}=0.5$ and $\Delta_{s}=0.2$.
Impact of BFSs on thermal Hall effects.—A defining characteristic of CTSCs is the quantized thermal Hall effect in the low-temperature limit. The quantized coefficient of this effect is directly proportional to the Chern number, expressed as $\kappa_{xy}=C\kappa_{0}$, where $\kappa_{0}=\pi k_{B}^{2}T/12\hbar$ [76, 77]. When a bulk energy gap exists, low-energy excitations are confined to the boundary, and the quantized effect can be derived using conformal field theory. This theory elucidates that each branch of the chiral Majorana edge state contributes equally to the effect [10]. However, in the presence of BFSs, low-energy excitations also emerge in the bulk. These low-energy Bogoliubov quasiparticles near the BFSs experience a nonzero Berry curvature, resulting in a nonzero contribution to the thermal Hall effect that is generally non-quantized [78]. To investigate the influence of BFSs on the thermal Hall effect, we utilize the formula for the thermal Hall coefficient developed in Refs. [79, 80], which is given by
$\kappa_{xy}=-\frac{k_{B}^{2}T}{2\hbar}\int_{-\infty}^{+\infty}dE\frac{E^{2}}{(k_{B}T)^{2}}\sigma(E)f^{\prime}(E),$ (12)
where $f(E)=1/[\exp(E/k_{B}T)+1]$ is the Fermi-Dirac distribution function, $k_{B}$ is the Boltzmann constant (we set $k_{B}=1$), $T$ is the temperature, and
$\sigma(E)=\sum_{n}\int_{E_{n}(\boldsymbol{k})<E}\frac{d^{2}k}{(2\pi)^{2}}\Omega^{(n)}(\boldsymbol{k}).$ (13)
@@
Figure 5: Evolution of the thermal Hall coefficient under the variation of $\eta$. The values of the shared parameters are: $t=1$, $\lambda_{\rm so}=0.5$, $M_{z}=0.5$ and $\Delta_{s}=0.2$. (a) $\mu=-4$, (b) $\mu=0$, (c) $\mu=1$, (d) $\mu=4$. The critical value of $\eta$ for the onset of BFSs is approximately equal to 0.11 in (a), 0.19 in (b), 0.27 in (c) and 0.72 in (d). The Chern number before the onset of BFSs is equal to 22 in (a), 4-4 in (b), 0 in (c) and $2$ in (d).
In Fig.5, we present the evolution of $\kappa_{xy}$ under the variation of $\eta$. Although the pairing amplitude depends on temperature, here we neglect this dependence. A detailed discussion of the temperature-dependent pairing case is provided in the SM [75]. We emphasize that this approximation is rather accurate across a wide range of low temperatures, which is the primary focus of our study.
@@
Specially, the Chern number in Figs.5(a), 5(b), 5(c) and 5(d) is $2$, 4-4, 0, and $2$, respectively, before $\eta$ surpasses the critical value, denoted as $\eta_{c}$, where BFSs emerge. Notably, $\eta_{c}$ depends on other parameters, as can be inferred from Eq.(11). The results depicted in Fig.5 demonstrate quantized behaviors in the low-temperature limit, provided that the bulk energy gap remains open. As anticipated, once $\eta$ exceeds $\eta_{c}$, the quantized behavior is disrupted due to the presence of BFSs. Intriguingly, we observe that the change in $\kappa_{xy}$ caused by BFSs is most pronounced in Fig.5(a). For the largest value of $\eta$ considered, $\kappa_{xy}$ in Fig.5(a) deviates significantly from its original quantized value, while the results in Fig.5(b) and Fig.5(d) show much smaller derivations. The remarkable change observed in Fig.5(a) can be attributed to the fact that the spectrum crossing of chiral Majorana edge states occurs at $k_{x}=0$ or $k_{y}=0$ for this particular case. As discussed earlier, the presence of BFSs causes these chiral Majorana edge states to disappear.
@@
Discussions and conclusions.—We demonstrated that applying a Zeeman field perpendicular to the spin plane of the inversion-symmetric SOC can transform the nonsymmorphic DSM into a Chern insulator. Furthermore, we showed that incorporating $s$-wave superconductivity into this system can lead to the realization of fully-gapped CTSCs with large Chern numbers, as well as gapless superconducting phases where BFSs coexist with chiral Majorana edge states. Additionally, we found that the presence of BFSs can have a significant impact on the thermal Hall effect.
In addition to the thermal Hall effect, the chiral Majorana edge states and BFSs can be experimentally probed using a scanning tunneling microscope to measure the density of states at the edges and in the bulk, respectively. The chiral Majorana edge states are characterized by a finite and constant density of states localized around the edges, while the BFSs contribute a finite density of states at zero energy across the system. Interestingly, we observe that in this system, the density of states develops a zero-bias peak at the onset of BFSs. This robust and distinctive signature is expected to be readily measurable [75]. Regarding the experimental realization of our proposed setup, we highlight that a variety of 2D DSMs have been theoretically predicted. These include black phosphorus [81], FeSe [82], X3SiTe6 ($X$ = Ta, Nb) [83, 84], SbSSn [85], HfGe0.92Te [86], and others [87], several of which have already been experimentally confirmed. Moreover, suitable superconductor substrates are abundant, as only conventional $s$-wave pairing is required. These factors ensure that our predictions are well within the realm of experimental feasibility.
In conclusion, our study underscores the rich interplay between Zeeman fields, superconductivity, and nonsymmorphic Dirac band structures, which can give rise to a diverse array of phases exhibiting intriguing properties.
@@
Acknowledgements.—This work is supported by the National Natural Science Foundation of China (Grant No.12174455), Natural Science Foundation of Guangdong Province (Grant No. 2021B1515020026), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023B1515040023).